# Attention Layers Add Into Low-Dimensional Residual Subspaces

### v1

@@ 5.3 Scaling Laws
To understand how our method scales, we evaluate performance as the number of SAE features increases from 1616K to 11M, keeping other hyperparameters fixed (details in Appendix E). Figure 7 summarizes the results.
@@
As shown in Figure 7(a), the normalized MSE follows a smooth power-law decay with increasing feature count. Across all scales, Active Subspace Init consistently outperforms baseline TopK and AuxK, achieving superior reconstruction at equivalent feature counts.
@@
Prior work (Bricken et al., 2023a) identified stale momentum as a key factor in dead feature formation. Building on this insight, we propose using SparseAdam, an optimizer specifically designed for sparse activation settings. By updating only the moments and parameters corresponding to non-zero gradients, SparseAdam naturally avoids stale momentum and thus mitigates the dead feature issue. As shown in Figures 7(a),  7(b), combining Active Subspace Init with SparseAdam substantially reduces dead features while reaching the lowest reconstruction error. While orthogonal to our initialization method, this choice provides a practical complement that further stabilizes training when scaling SAEs to very large capacities. We discuss more about stale momentum and SparseAdam in Appendix A.
@@ 5.4 Generalize to Sparse Replacement Models
To test this, we apply Active Subspace Initialization to Lorsa [^5] Specifically, we initialize the feature directions (corresponding to $W^{O}$ in Lorsa) within the active subspace. Other hyperparameters are the same as He et al. (2025).. This modification reduces the proportion of dead parameters from $68.4\%$ to $40.5\%$ under identical hyperparameter settings (Table 1), while also slightly improving reconstruction error.
Table 1: Effect of Active Subspace Initialization on reducing dead parameters in attention replacement model.
Method Vanilla Active Subspace Init
Dead Parameters (%) 68.4 40.5
Normalized MSE 0.130 0.121
These results indicate that Active Subspace Initialization provides an effective strategy for mitigating dead parameters when training sparse replacement models on low-rank activations. We posit that complete elimination of dead heads may require additional mechanisms, such as the tied initialization used in SAEs to ensure alignment between feature encoding and decoding methods666”Match” means encoder can be initialized to predict relatively accurate feature activation values for decoder.. This approach has been shown to be crucial for reducing dead features in SAEs (Gao et al., 2024), and its absence in Lorsa may limit further improvement.
@@ 6 Conclusion
Appendix A Stale Momentum as Another Root Cause of Dead Features
Recent work by Bricken et al. [2023a] identifies stale momentum as a key cause to dead feature formation. Specifically, when a feature remains inactive over training steps, its associated optimizer momentum continues to accumulate. If the feature activates, the stale momentum results in disproportionately large updates, destabilizing training and potentially suppressing that feature permanently.
@@ Appendix A Stale Momentum as Another Root Cause of Dead Features
Appendix B Activation Sources
@@
Unless otherwise specified, activations are extracted from intermediate layers. For example, in LLaMA-3.1-8B (32 layers), we use activations from layer 15 (zero-indexed). We analyze four types of activations: (1) attention output, (2) MLP output, and (3) residual stream (post layer). For each type, we collect 10 million activation vectors. We empirically verify that this sample size is sufficient to produce stable and reproducible spectral analyses.
Appendix C More Low-Rank Result Across Different Models and Datasets
@@
We present relative singular values and fraction of loss recovered for some other model-dataset pairs in Figure 8. Models include pythia-2.8b [^10] EleutherAI/pythia-2.8b. Datasets include RedPajamaGithub [^11] https://huggingface.co/datasets/cerebras/SlimPajama-627B and CCI3-Data [^12] https://huggingface.co/datasets/BAAI/CCI3-Data
@@
Appendix D Different Choose of Variance Threshold for Intrinsic Dimension
@@
Appendix E SAE Training Details
E.1 Collecting Activations
@@
E.2 Initialization and Optimization
The decoder columns $W^{dec}_{:,i}$ are initialized uniformly and normalized to achieve the lowest initial reconstruction loss. We find that the specific initialization norm has little impact, as long as in a reasonable scope. For example, initializing $W^{dec}_{:,i}$ uniformly with a fixed bound, as in Conerly et al. [2025], yields similar results. The encoder weights $W^{enc}$ are initialized as the transpose of $W^{dec}$, while both the encoder bias $b^{enc}$ and decoder bias $b^{dec}$ are set to zero.
This initialization scheme ensures that the SAE begins training with an almost zero reconstruction loss. Such initialization has been widely observed to benefit SAE training.
We train SAEs using the Adam and SparseAdam optimizers, both with $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=10^{-8}$.
E.3 Fixed Hyperparameters in Scaling Law
@@
We empirically set $batchsize=4096$, which belows the critical batch size.
@@
The learning rate for Adam and SparseAdam is sweeped separately in $1\mathrm{e}{-5}$, $2\mathrm{e}{-5}$, $4\mathrm{e}{-5}$, $6\mathrm{e}{-5}$, $8\mathrm{e}{-5}$, $1\mathrm{e}{-4}$, $2\mathrm{e}{-4}$, $4\mathrm{e}{-4}$], and we ultimately use $4\mathrm{e}{-5}$ for Adam and $6\mathrm{e}{-5}$ for SparseAdam.
@@
We follow Gao et al. [2024] to set auxiliary loss coefficient $\alpha$ as $\frac{1}{32}$. We sweep the $k_{aux}$ in [256, 512, 1024, 2048] and finally choose 512.
@@
As shown in Figure 5, $d_{init}$ is a hyperparameter with a wide range of sub-optimal value space (from 256 to 2048). We use 768 for all experiments.
@@
We use $2.5B$ tokens for each SAE training.

### v2

@@ 5.3 Scaling Laws
To understand how our method scales, we evaluate performance as the number of SAE features increases from 1616K to 11M, keeping other hyperparameters fixed (details in Appendix G.5).
@@
As shown in Figure 7(a), Active Subspace Init consistently outperforms TopK and AuxK across all scales.
@@
Prior work (Bricken et al., 2023a) identified stale momentum as a key factor in dead feature formation. Building on this insight, we propose using SparseAdam, an optimizer specifically designed for sparse activation settings. By updating only the moments and parameters corresponding to non-zero gradients, SparseAdam naturally avoids stale momentum and thus mitigates the dead feature issue. As shown in Figures 7(a),  7(b), combining Active Subspace Init with SparseAdam substantially reduces dead features while reaching the lowest reconstruction error. While orthogonal to our initialization method, this choice provides a practical complement that further stabilizes training when scaling SAEs to very large capacities. We discuss more about stale momentum and SparseAdam in Appendix C.
@@ 5.4 Generalize to Sparse Replacement Models
To evaluate this, we apply Active Subspace Initialization to Lorsa. This modification reduces the proportion of dead parameters significantly under identical hyperparameter settings (Figure 8), while also improving the reconstruction. Specifically, we initialize the decoder (corresponding to $W^{O}$ in Lorsa) within the active subspace of each original MHSA head, while the encoder (corresponding to $W^{V}$ in Lorsa) is aligned with the corresponding input-side directions of the decoder. Further implementation details are provided in Appendix I.
Figure 8: Effect of Active Subspace Initialization on reducing dead parameters in attention replacement model.
6 Discussion
When to Use Active Subspace Initialization
For activation sites that do not exhibit clear low-rank structures (e.g., residual stream), our method yields only limited gains (Appendix B.2). The singular value spectrum provides a more reliable indicator: when the spectrum decays rapidly in its tail, applying ASI is likely to be effective.
Causality between Low-Rank Structure and Dead Features
We observed a strong correlation between low-rank structure and the emergence of dead features (Section 5), but the causal mechanism remains unclear. It may be tied to optimization dynamics or feature competition, yet a precise explanation is lacking. A deeper understanding of this relationship could also explain why ASI is less effective under high-$\ell_{0}$ warm-up regimes (Appendix H). We leave this important line of inquiry to future work.
@@ 7 Conclusion
Appendix A Declaration of LLM Usage
We acknowledge the use of large language models (LLMs) to assist in the preparation of this manuscript. Specifically, ChatGPT and Claude were utilized during the writing and coding process for the exclusive purpose of improving text clarity, grammar, overall fluency, and for generating boilerplate code structures. All ideas, theoretical developments, experimental designs, results, analyses, and scientific conclusions are entirely our own. The LLMs acted solely as assistive tools and were not involved in any aspect of the intellectual or scientific work.
Appendix B Ablation Study
B.1 Active Subspace Init vs Random Subspace Init
We employ random subspace initialization as a baseline and observe that it consistently degrades SAE training across all metrics, as shown in Figure 9.
Figure 9: For activations with a full space dimension of 4096, proportion of dead features (left), normalized MSE (mid) and Delta LM loss (right) across different subspace dimensions. Random subspaces are used as the baseline, whereas only initialization with the active subspace yields improvement.
B.2 Apply on Near-Full-Rank Activation
We also apply Active Subspace Initialization (ASI) to near-full-rank activations, such as those in the residual stream, to evaluate its generality. When training an SAE on the post-layer-15 residual stream of Llama-3.1-8B, we find ASI yields minimal gains (Figure 10). This is consistent with our expectation, as these activations inherently exhibit a lower rate of dead features even with standard initialization.
Figure 10: For activations with a full space dimension of 4096, proportion of dead features (left), normalized MSE (mid) and Delta LM loss (right) across different subspace dimensions. Random subspaces are used as the baseline, whereas only initialization with the active subspace yields improvement.
B.3 Choice of the Initial Dictionary Size $D_{init}$
As shown in Figure 5, $D_{init}$ is a hyperparameter with a wide range of acceptable values (from 256 to 2048). We hypothesize that the performance degradation at very low values occurs because an excessively small dictionary fails to span the subspace containing critical information needed for effective reconstruction.
Appendix C Stale Momentum as Another Root Cause of Dead Features
Recent work by Bricken et al. (2023a) identifies stale momentum as a key cause to dead feature formation. Specifically, when a feature remains inactive over training steps, its associated optimizer momentum continues to accumulate. If the feature activates, the stale momentum results in disproportionately large updates, destabilizing training and potentially suppressing that feature permanently.
@@ Appendix C Stale Momentum as Another Root Cause of Dead Features
Appendix D Activation Sources
@@
Unless otherwise specified, activations are extracted from model intermediate layers. For example, in LLaMA-3.1-8B (32 layers), we use activations from layer 15 (zero-indexed). We analyze three types of activations: (1) attention output, (2) MLP output, and (3) residual stream (post layer).
Appendix E More Low-Rank Result Across Different Models and Datasets
@@
We present relative singular values and fraction of loss recovered for some other model-dataset pairs in Figure 11. Models include pythia-2.8b [^11] https://huggingface.co/EleutherAI/pythia-2.8b. Datasets include RedPajamaGithub [^12] https://huggingface.co/datasets/cerebras/SlimPajama-627B and CCI3-Data [^13] https://huggingface.co/datasets/BAAI/CCI3-Data
@@
Appendix F Different Choose of Variance Threshold for Intrinsic Dimension
@@
Appendix G SAE Training Details
G.1 Collecting Activations
@@
G.2 Initialization
The decoder columns $W^{dec}_{:,i}$ are initialized uniformly, and the optimal norm for them is found through a grid search to minimize the initial reconstruction loss. We find that the specific initialization norm has little impact, as long as in a reasonable scope. For example, initializing $W^{dec}_{:,i}$ uniformly with a fixed bound, as in Conerly et al. (2025), yields similar results. The encoder weights $W^{enc}$ are initialized as the transpose of $W^{dec}$, while both the encoder bias $b^{enc}$ and decoder bias $b^{dec}$ are set to zero.
G.3 Optimization
We train SAEs using the Adam optimizers, both with $lr=2*10^{-4}$, $batchsize=32768$, $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=10^{-8}$. The differences in scaling law experiments are described in the Appendix G.5.
G.4 Jumprelu SAEs
We trained JumpReLU SAEs (Rajamanoharan et al., 2024) under two distinct hyperparameter configurations: one maintaining a consistently low $\ell_{0}$ value throughout training, and another where $\ell_{0}$ is gradually decreased from a higher initial value. Unless otherwise specified, all JumpReLU SAEs were trained using the same settings as Conerly et al. (2025), which corresponds to the latter configuration. The key modifications for the former setting are as follows: (1) we initialized the encoder bias to zero instead of applying the heuristic that equalizes feature activation counts at initialization, and (2) we kept the sparsity coefficient fixed rather than employing a global warm-up schedule. As a result, the $\ell_{0}$ sparsity level started at a relatively low value early in training. This design is critical to our approach: we observed that if the model remains in a high-$\ell_{0}$ regime (e.g., on the order of $d_{\text{model}}/2$) for an extended period before sparsity increases, the feature directions tend to drift away from the active subspace during this phase, thereby diminishing the effectiveness of our method (Appendix H.1).
G.5 Fixed Hyperparameters in Scaling Law
@@
We empirically set the batch size to 40964096—below the critical batch size—which differs from the setting of $32768$ used in other experiments in this paper. This reduction is necessary because training SAEs with 1 million features requires substantial memory, posing significant computational challenges.
@@
The learning rate for Adam and SparseAdam is sweeped separately in $1\mathrm{e}{-5}$, $2\mathrm{e}{-5}$, $4\mathrm{e}{-5}$, $6\mathrm{e}{-5}$, $8\mathrm{e}{-5}$, $1\mathrm{e}{-4}$, $2\mathrm{e}{-4}$, $4\mathrm{e}{-4}$], and we ultimately use $4\mathrm{e}{-5}$ for Adam and $6\mathrm{e}{-5}$ for SparseAdam. We employ a three-phase learning rate schedule consisting of a linear warm-up, a stable phase, and a linear decay. The learning rate increases linearly from zero to its maximum value over the first 500 steps, remains constant during the intermediate phase, and then decays linearly to 1% of the maximum value over the final 20% of the total training steps.
@@
We follow Gao et al. (2024) to set auxiliary loss coefficient $\alpha$ as $\frac{1}{32}$. We sweep the $k_{aux}$ in [256, 512, 1024, 2048] and finally choose 512.
@@
We use 768 for all experiments.
@@
We use 2.52.5B tokens for each SAE training.
Appendix H Use ASI on other Activation Functions
H.1 Jumprelu
Another wildly used activation fuction is Jumprelu (Rajamanoharan et al., 2024). We trained the Jumprelu SAEs under two different hyperparameter settings: one with a consistently low $\ell_{0}$ value and another where $\ell_{0}$ gradually decreases from a higher initial value, as described in Appendix G.4. We observed that our method is effective in the former case (Figure 13) but shows little improvement in the latter (Figure 14).
For cases where one follows a schedule that gradually reduces $\ell_{0}$ from a high initial value, we recommend first applying PCA to reduce the dimensionality of the data. The SAE can then be trained on the reduced representation until the $\ell_{0}$ level reaches the target range. Afterwards, the PCA projection matrix can be folded into the model parameters, and training can continue in the original space. This achieves a similar effect without the drawbacks of prolonged training in the high-$\ell_{0}$ regime.
Figure 13:
Figure 14:
H.2 TopK with K Anneal
To enhance the finding in Section H.1, we conduct experiments on a variant of TopK, which sets K to a high value and then lets it decrease during training (He et al., 2024). We find ASI also fails in this case (Figure 15).
Figure 15:
Appendix I Pseudo-code for implementing Active Subspace Init
Below is a PyTorch-style pseudo-code for Active Subspace Initialization.
Use on SAE
# X: activation batch [batch_size, d_model]
# W_E: decoder weight [d_model, d_sae]
# W_D: decoder weight [d_sae, d_model], initialized uniformly
# d_active_subspace: target subspace dimension
# 1. Demean the activations
demeaned_X = X - X.mean(dim=0) # [batch_size, d_model]
# 2. Compute SVD
U, S, V = torch.svd(demeaned_label) # V: [d_model, d_model]
# 3. Take top-d_init singular vectors
proj_weight = V[:, :d_init]  # [d_model, d_init]
# 4. Fold projection into decoder weights
W_D.copy_(W_D[:, :d_init] @ proj_weight.T)
# 5. Init W_E with W_D.T
W_E.copy_(W_D.T)
Use on Lorsa
# X: activation batch [batch_size, seq_len, d_model]
# mhsa: pretrained MHSA module with W_O and W_V
# W_O, W_V: Lorsa decoder/encoder weights
# d_head: dimension of original MHSA heads
# n_heads: number of original MHSA heads
# n_lorsa_heads: number of Lorsa heads
# 1. Run MHSA to capture per-head outputs
Z = mhsa.compute_z(X) # [batch_size, seq_len, n_heads, d_head]
output_per_head = torch.einsum(’bsnh, nhd->bsnd’, Z, mhsa.W_O)
# 2. For each original head
# initialized some lorsa heads into it’s active subspace
rate = n_lorsa_heads // n_heads
for mhsa_index in range(n_heads):
    head_slice = [rate*mhsa_index:rate*(mhsa_index+1)]
    # [B, S, d_model]
    output=output_per_head[:, :, mhsa_index, :]
    output_flat=output.flatten(0, 1) # [B*S,d_model]
    # 3.1 Demean
    demeaned_output=output_flat - output_flat.mean(dim=0)
    # 3.2 Compute SVD
    U, S, V = torch.svd(demeaned_output)
    # 3.3 Take top-d_head singular vectors
    proj_weight = V[:, :d_head]
    # 3.4 Update part of decoder weights W_O
    W_O[head_slice] = W_O[head_slice, :d_head] @ proj_weight.T
    # 3.5 Update part of encoder weights W_V
    head_trans = mhsa.W_V[mhsa_index] @ mhsa.W_O[mhsa_index]
    W_V[head_slice] = W_O[head_slice] @ head_trans.T
The strategy of initialize $W_{V}$ in Lorsa is a method like the tied initialization used in SAEs to ensure alignment between feature encoding and decoding141414”Match” means encoder can be initialized to predict relatively accurate feature activation values for decoder.. This approach has been shown to be crucial for reducing dead features in SAEs (Gao et al., 2024). We think the same thought could also be used to improve the replacement model for MLP (trancoder and cross layer transcoder), which we leave a deeper investigation to future work.